OpenStudy (noahbred):

HELPPP? If an object with a mass of 6 kg is pushed with a force of 42 N, at what rate will this object accelerate?

OpenStudy (noahbred):

@Abhisar

OpenStudy (abhisar):

Force = Mass \(\times\) Acceleration Use this formula

OpenStudy (noahbred):

so 6x42=242

OpenStudy (abhisar):

42 = 6 \(\times\) acceleration

OpenStudy (abhisar):

So acceleration will be how much?

OpenStudy (noahbred):

IS IT 7 ?!?!

OpenStudy (noahbred):

or do i divide

OpenStudy (abhisar):

Yes, it's 7

OpenStudy (shamim):

Given Acceleration=? Force=42 N Mass=6kg

OpenStudy (anonymous):

treat object as particle. Newtonian mechanics condition.

OpenStudy (anonymous):

f=ma a=f/m a=42/6 a=7 thats it its soo simple